# Metropolis-Hastings Sampling for 3D Gaussian Reconstruction

### v1

@@ Appendix C Extended Derivation for Surrogate
This section explains why we approximate $-\Delta\mathcal{L}$ to $I(i)$.
@@
where $\nabla_{g_{i}}\,\mathcal{L}(\Theta)$ is the gradient of the loss w.r.t the Gaussian’s parameters and the inner product measures the first-order change when adding $\varepsilon\,g_{i}$.
@@
$-\,\Delta\mathcal{L}\;\propto\;\alpha\,O(p_{i})\;+\;\beta\,\mathrm{SSIM}_{\rm agg}(p_{i})\;+\;\gamma\,\mathcal{L}1_{\rm agg}(p_{i})\;=\;z_{i}$ (32)
As stated in Eq. 10 and 11, we set:
@@
$-\,\Delta\mathcal{L}\approx z_{i}\approx\sigma(z_{i})=I(i)$ (34)
@@ Appendix D Algorithms
Our algorithms of our coarse-fine Metropolis-Hastings sampling framework and overall pipeline are summarized in Algorithm 1 and 2.
@@
2:Camera set $\mathcal{C}$; iteration $t$, total steps $t_{\max}$
@@
6:// Select diverse camera subset for error aggregation
7:$\mathcal{C}_{t}\leftarrow\textsc{CameraSubset}(\mathcal{C},t,t_{\max})$
@@
10:$\mathrm{SSIM}_{\text{agg}},\mathcal{L}1_{\text{agg}}\leftarrow\frac{1}{|\mathcal{C}_{t}|}\sum_{c\in\mathcal{C}_{t}}\left[1{-}\mathrm{SSIM}(I^{(c)},\hat{I}^{(c)}),\;\;|I^{(c)}-\hat{I}^{(c)}|\right]$
@@
19:    $p_{i}\leftarrow\Pi(\mathbf{x}_{i})$triangleright Project center to image plane
20:    $O(p_{i})\leftarrow\alpha_{i}$triangleright Use projected Gaussian opacity
@@
33:        $\rho\leftarrow I(i)\cdot D(v^{\prime})$triangleright Acceptance probability
@@
6:     $I\leftarrow{\footnotesize\textsc{Rasterize}}(M,S,C,A,V)$
7:     $L\leftarrow{\footnotesize\textsc{Loss}}(I,\hat{I})$
8:     $S_{\text{SSIM},i}\leftarrow{\footnotesize\textsc{SSIM}}(I,\hat{I})$
9:     $S_{\text{L1},i}\leftarrow\text{L1}(I,\hat{I})$
10:     $[M,S,C,A]\leftarrow{\footnotesize\textsc{Adam}}(\nabla L)$triangleright Update attributes
@@
Appendix E Comparison of the Number of Gaussians
Table 4 contrasts the number of Gaussians needed by our approach and by 3DGS for each scene. In every case, our method attains equal or superior rendering quality while requiring substantially fewer Gaussians.
Table 4: Comparison of our method and 3DGS on the number of Gaussian center points used to represent scenes in the Mip-NeRF 360 [1], Tanks & Temples [21], and Deep Blending [15] datasets.
Dataset Mip-NeRF 360 Tanks&Temples Deep Blending
Appendix F Preliminary on Metropolis-Hastings Algorithm
The Metropolis-Hastings algorithm [33, 7] is a widely used Markov Chain Monte Carlo (MCMC) method designed to sample from complex, high-dimensional probability distributions. The objective of the algorithm is to construct a Markov chain whose stationary distribution $\pi(\mathbf{x})$ matches the target distribution, enabling efficient sampling where direct methods are impractical.
Given a target distribution $\pi(\mathbf{x})$ defined over a state space $\mathcal{X}$, the Metropolis-Hastings algorithm generates a sequence $\{\mathbf{x}^{(t)}\}_{t=1}^{N}$ of samples by iteratively proposing candidate points and accepting or rejecting them based on an acceptance probability. The transition from the current state $\mathbf{x}^{(t)}$ to a new candidate $\mathbf{y}$ is governed by a proposal distribution $q(\mathbf{x},\mathbf{y})$.
At each iteration $t$, a new candidate $\mathbf{y}$ is proposed from the distribution $q(\mathbf{x}^{(t)},\cdot)$. The acceptance probability is further computed as
Appendix G Per-Scene Results
In this section, we provide the full per-scene evaluations. We evaluated our method, 3DGS [18] and 3DGS-MCMC [20] on various datasets and scenes. We provide the per-scene results for Mip-NeRF 360 [1], Tanks & Temples [21] and Deep Blending [15] datasets in Table 5 and 6. All evaluations have been conducted on a NVIDIA GeForce RTX 3090.
Table 5: Full per-scene results evaluated on Mip-NeRF 360 dataset.
@@ Appendix G Per-Scene Results
Table 6: Full per-scene results evaluated on Tank&Temples and Deep Blending datasets.
@@
3DGS-MCMC PSNR ($\uparrow$) 22.58 26.02 24.30 29.37 30.30 29.84
@@
Appendix H Comprehensive Ablation Study for Relocation
This section reports the full results of our relocation ablation study. We provide the results on the Mip-NeRF 360 [1], Tanks & Temples [21], and Deep Blending [15] datasets in Table 7.
Table 7: Ablation on relocation. Results are presented for both 3DGS and 3DGS-S, the latter using the same number of Gaussians as ours.
Mip-NeRF360 [1] Tanks&Temples [21] Deep Blending [15]
[…]
3DGS [18] 27.44 / 0.811 / 0.223 23.63 / 0.848 / 0.177 29.53 / 0.904 / 0.244
3DGS-S [18] 26.62 / 0.761 / 0.299 23.38 / 0.828 / 0.214 29.27 / 0.897 / 0.272
Ours w/o Reloc. 26.94 / 0.778 / 0.267 23.69 / 0.844 / 0.184 29.57 / 0.901 / 0.267
Ours 27.34 / 0.798 / 0.241 23.99 / 0.852 / 0.166 30.12 / 0.909 / 0.245
Appendix I Comprehensive Ablation Study for Importance Score Components
This section reports the full results of our ablation study on the importance score components. We provide the results on the Mip-NeRF 360 [1], Tanks & Temples [21], and Deep Blending [15] datasets in Table 8.
Table 8: Full Ablation Study Results on Importance Score Components.
Method | Dataset Mip-NeRF360 [1] Tanks&Temples [21] Deep Blending [15]
L1 SSIM Opacity PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
27.22 0.796 0.243 23.89 0.851 0.168 29.87 0.907 0.247
27.30 0.796 0.242 23.92 0.851 0.169 29.85 0.907 0.246
27.30 0.796 0.243 23.87 0.852 0.167 30.02 0.908 0.247
27.34 0.798 0.241 23.99 0.852 0.166 30.12 0.909 0.245
Appendix J Broader Impacts
@@ Appendix I Comprehensive Ablation Study for Importance Score Components
Appendix K Dataset Licenses

### v2

@@ Appendix C Extended Derivation for Surrogate
This section explains why we approximate $\Delta\mathcal{L}$ to $I(i)$.
@@
where $\nabla_{g_{i}}\,\mathcal{L}(\Theta)$ is the gradient of the loss w.r.t. the Gaussian’s parameters and the inner product measures the first-order change when adding $\varepsilon\,g_{i}$.
@@
$\Delta\mathcal{L}\;\propto\;\alpha\,O(p_{i})\;+\;\beta\,\mathrm{SSIM}_{\rm agg}(p_{i})\;+\;\gamma\,\mathcal{L}1_{\rm agg}(p_{i})\;=\;z_{i}$ (34)
As stated in Eq. 12 and 13, we set:
@@
$\Delta\mathcal{L}\approx z_{i}\approx\sigma(z_{i})=I(i)$ (36)
@@ Appendix D Algorithms
Our algorithms for our coarse-fine Metropolis-Hastings sampling framework and overall pipeline are summarized in Alg. 1 and 2.
@@
2:View set $\mathcal{C}$; iteration $t$, total steps $t_{\max}$
@@
6:// Select diverse view subset for error aggregation
[…]
@@
[…]
@@
19:  $p_{i}\leftarrow\Pi(\mathbf{x}_{i})$ $\triangleright$ Project center to image plane
20:  $O(p_{i})\leftarrow\alpha_{i}$ $\triangleright$ Use projected Gaussian opacity
@@
33:   $\rho\leftarrow I(i)\cdot D(v^{\prime})$ $\triangleright$ Acceptance probability
@@
5:  $\{V,\hat{I}\}\leftarrow{\footnotesize\textsc{SampleTrainingView}}()$ $\triangleright$ Camera $V$ and image
6:  $I\leftarrow{\footnotesize\textsc{Rasterize}}(M,S,C,A,V)$
@@
[…]
10:  $M,S,C,A]\leftarrow{\footnotesize\textsc{Adam}}(\nabla L) \triangleright$ Update attributes
11:  if ${\footnotesize\textsc{IsRefinementIteration}}(i)$ then
@@
Appendix E Per-Scene Results
In this section, we provide the full per-scene evaluations. We evaluated our method, 3DGS [20] and 3DGS-MCMC [22] on various datasets and scenes. We provide the per-scene results for Mip-NeRF 360 [1], Tanks & Temples [23] and Deep Blending [17] datasets in Tab. A and B. All evaluations have been conducted on a NVIDIA GeForce RTX 3090.
Table A: Full per-scene results evaluated on Mip-NeRF 360 dataset.
@@ Appendix E Per-Scene Results
Table B: Full per-scene results evaluated on Tanks & Temples and Deep Blending datasets.
@@
3DGS-MCMC [22] PSNR ($\uparrow$) 22.58 26.02 24.30 29.37 30.30 29.84
@@
Appendix F Comparison of the Number of Gaussians
Table C contrasts the number of Gaussians needed by our approach and by 3DGS for each scene. In every case, our method attains equal or superior rendering quality while requiring substantially fewer Gaussians.
Table C: Comparison of our method and 3DGS on the number of Gaussian center points used to represent scenes in the Mip-NeRF 360 [1], Tanks & Temples [23], and Deep Blending [17] datasets.
Dataset Mip-NeRF 360 Tanks&Temples Deep Blending
Appendix G Broader Impacts
@@
Appendix H Dataset Licenses